INFRARED SPECTRA OF COMPLEXES CONTAINING ACETYLENE-d2

CLÉMENT LAUZIN, J. NOROOZ OLIAEE, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, 2500 University Dr., N.W., Calgary, AB T2N 1N4, Canada; \underline{A.R.W. MCKELLAR}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, ON K1A 0R6, Canada.

Infrared spectra of the C_{2}D_{2} dimer in the monomer ν_{3} region (∼2439 cm^{−1}) are observed by direct absorption using a rapid-scan tuneable diode laser spectrometer to probe a pulsed supersonic slit-jet expansion. We analyze the perpendicular K = 1-0 and 0-1 subbands of the vibrational mode involving the C_{2}D_{2} monomer located at the top of the T-shaped dimer, but miss the parallel band involving the stem monomer vibration due to limited laser coverage. The results are consistent with previous work on acetylene dimers, but perturbations are much less evident than in the analogous infrared spectrum of C_{2}H_{2}.^{a} As expected, the tunneling splitting in the excited state (8 MHz) is much smaller than in the ground state (424 MHz).^{b} In the same region, we observe the H-bonded isomer of the C_{2}D_{2}-C_{2}H_{2} dimer. This has not previously been observed, even though microwave spectra of almost every other conceivable deuterated isotopologue are known.^{b} In addition to these acetylene dimers, our spectra also contain bands arising from impurities in the gas mixture which we assign to the C_{2}D_{2}-nitrogen and C_{2}D_{2}-water complexes

aG.T. Fraser, R.D. Suenram, F.J. Lovas, A.S. Pine, J.T. Hougen, W.J. Lafferty, and J.S. Muenter, J. Chem. Phys. 89, 6028 (1988).

bK. Matsumura, F.J. Lovas, and R.D. Suenram, J. Mol. Spectrosc. 150, 576 (1991).